# Euclid Quick Data Release (Q1)

### v2

@@
Light emission from galaxies exhibit diverse brightness profiles, influenced by factors such as galaxy type, structural features and interactions with other galaxies. Elliptical galaxies feature more uniform light distributions, while spiral and irregular galaxies have complex, varied light profiles due to their structural heterogeneity and star-forming activity. In addition, galaxies with an active galactic nucleus (AGN) feature intense, concentrated emission from gas accretion around supermassive black holes, superimposed on regular galactic light, while quasi-stellar objects (QSO) are the extreme case of the AGN emission dominating the galaxy. The challenge of identifying AGN and QSO has been discussed many times in the literature, often requiring multi-wavelength observations. This paper introduces a novel approach to identify AGN and QSO from a single image. Diffusion models have been recently developed in the machine-learning literature to generate realistic-looking images of everyday objects. Utilising the spatial resolving power of the Euclid VIS images, we created a diffusion model trained on one million sources, without using any source pre-selection or labels. The model learns to reconstruct light distributions of normal galaxies, since the population is dominated by them. We condition the prediction of the central light distribution by masking the central few pixels of each source and reconstruct the light according to the diffusion model. We further use this prediction to identify sources that deviate from this profile by examining the reconstruction error of the few central pixels regenerated in each source’s core. Our approach, solely using VIS imaging, features high completeness compared to traditional methods of AGN and QSO selection, including optical, near-infrared, mid-infrared, and X-rays. Our study offers practical insights for refining diffusion models and broadening their applications throughout the Euclid survey area, underscoring the utility of this approach in diverse astronomical contexts beyond just AGN identification.
@@ 1 Introduction
Ever since the discovery that the mass or dynamics of a spheroid within a galaxy correlates with the mass of any detectable supermassive black hole (SMBH) at its centre (e.g., Dressler 1989; Kormendy & Richstone 1995; Magorrian et al. 1998; Ferrarese & Merritt 2000, and more recently, Sahu et al. 2019; Davis et al. 2018, 2019 and references therein), it has been believed that most – if not all – galaxies with a spheroidal (or pseudo-spheroidal) component contain a SMBH at their centre. If a galaxy’s emission shows evidence for its central SMBH being actively fuelled, it will emit non-thermal radiation, leading it to be classified as an active galactic nucleus (AGN). Depending upon the luminosity of that emission and its contrast with the stellar emission from the surrounding galaxy, it could have a classification ranging from a Seyfert galaxy (for the weakest emission) through to a quasi-stellar object or QSO (for the most luminous).
@@
For imaging in wavebands such as the optical and near-IR, the presence and detectability of emission from the area directly around a SMBH depends on multiple factors. This can include the rate of infall of material into the region around the SMBH, the amount of obscuration (absorption or scattering) towards that region, the contrast between the SMBH-related emission and that of the surrounding galaxy, the redshift, and the intrinsic compactness of the stellar distribution of the host galaxy. The growth and mass of a central SMBH is intimately related to the evolution of its host galaxy, not least through the action of energy, momentum and radiative feedback from outflows originating from the area around the SMBH on the properties of the stellar populations and ISM (see Harrison & Ramos Almeida 2024 for a recent review). Consequently, identifying and characterising SMBHs, particularly during periods of active feedback, is central to the study of SMBH and galaxy evolution.
@@
A more precise classification of a source as an AGN is possible when a much wider range of multi-wavelength data is available. This can include spectroscopy and considering the much broader spectral energy distribution of the galaxy and/or the candidate AGN. However, it is still worth considering the efficient selection of AGN candidates from single-band imaging as a first step in identifying large samples of AGN for further study. In particular, with the advent of the uniform, deep and high spatial-resolution optical imaging data set from the Euclid mission, covering a large fraction of the extragalactic sky, there is the potential to derive unprecedentedly large samples of AGN across a wide range of (optical) luminosity and redshift parameter space with the use of an appropriate and sufficiently efficient technique. The technique can then be subsequently refined by feeding back the result of using other multi-wavelength data, allowing us to confirm or reject candidates. With sufficient refinement, this further assessment step eventually becomes either unnecessary, or at the very least, an efficient stage in compiling reliable samples. This then opens the door to the production of unprecedentedly large and reliable samples of optically-selected AGN.
@@
Figure 1: The diffusion pipeline (top) progressively adds noise to images, training the model to predict what noise was added from the previous step. Once trained and during inference, the model takes pure Gaussian noise as input and can iteratively remove the noise until a realistic galaxy image remains. Repeat inference runs will provide a different and unique galaxy from those it was trained on. The Repaint pipeline (bottom) takes the trained diffusion model and enables conditioning to allow parts of an existing image to be preserved by masking. At each denoising step, noise levels in the preserved pixels are adjusted to ensure they integrate correctly with the newly generated sections. After $T$ iterations, the output includes the retained pixels and newly generated areas, creating a different yet plausible final image.
@@
This work presents a technique that can achieve the above goal. In the following, we describe the use of diffusion models for generating samples of AGN candidates using Euclid VIS $I_{\scriptscriptstyle\rm E}$ band (Euclid Collaboration: Mellier et al. 2025; Euclid Collaboration: Cropper et al. 2025) imaging alone. Diffusion-based methods are machine-learning models that generate or reconstruct data by evolving a random distribution of pixels into a structured output over several steps. This technique outperforms traditional methods like median pixel substitution or interpolation, which often struggle to recreate complex structures within images. Using these models for ‘inpainting’, where we adaptively recreate parts of the image, we can measure the discrepancies between the original and generated pixels. Examining these errors allows us to separate AGN, QSOs, and other sources from the broader galaxy population.
@@
Our method identifies potential active galactic nuclei (AGN) by treating the task as an unsupervised anomaly detection problem. It begins with an image of a galaxy, potentially containing AGN emission at its center, and masks out the central pixels that might be influenced by such emission. A diffusion model, trained on a large, unlabeled data set of galaxies – most of which are normal, non-active – then predicts or ‘inpaints’ the masked region based on the learned prior distribution of typical galaxy morphologies. Because AGN are rare, there is no requirement to remove them from the training sample; the model is inherently biased toward reconstructing the central region as if it belonged to a typical, non-active galaxy. As a result, any significant discrepancy between the inpainted and actual central region suggests the presence of an AGN. This reconstruction error thus serves as a powerful anomaly score for identifying candidate AGN.
@@
While this paper details the fundamental technique applied to VIS imaging data, as this approach utilises machine learning, future improvements could incorporate findings from a detailed assessment that involves additional data sets. These could include multi-wavelength cross-matching or spectroscopic validation, from which we can ultimately reduce the need for further time-expensive refinement steps. Such enhancements would help to further improve the accuracy and reliability of the AGN candidate samples generated by the technique. Additionally, the technique as presented here uses a standard treatment of diffusion models, which in principle can be significantly further optimised for the specific characteristics of the Euclid data set, thereby improving further the efficiency of the processing of the Euclid data. Such work would be the subject of future papers.
@@ 2.1 Generative models
Generative models are a class of machine learning systems designed to learn an underlying data distribution, $p(x)$, from a finite set of independent and identically distributed (i.i.d.) samples $\{x_{i}\}^{N}_{i=1}\sim p(x)$. The primary goal of these models is to develop a tractable method for generating new, synthetic samples that are statistically indistinguishable from the training data. Over the years, several distinct families of generative models have emerged, each with unique architectural and training paradigms. A key distinction lies in their optimisation objectives, which often seek to minimise a measure of divergence between the true data distribution $p(x)$ and the model’s learned distribution $p_{\theta}(x)$. The choice of divergence has significant implications for model behavior. Prominent examples include Variational Autoencoders (VAEs, Kingma 2013), which learn a latent variable model of the data, and Generative Adversarial Networks (GANs, Goodfellow et al. 2014, 2020), which involve a contest between a generator and a discriminator. GANs implicitly optimise an objective related to the reverse Kullback-Leibler (KL) divergence, $D_{KL}(p_{\theta}(x)||p(x))$, which exhibits a ‘mode-seeking’ property. This encourages the generation of high-fidelity samples but can lead to mode collapse, where the model learns only a subset of the data distribution (Thanh-Tung & Tran 2020). In contrast, VAEs and Diffusion Models optimise objectives related to the forward KL divergence, $D_{KL}(p(x)||p_{\theta}(x))$, which is ‘mass-covering’. This property encourages the model to account for all modes present in the data, leading to greater sample diversity and more comprehensive coverage of the data distribution.
@@ 2.2 Diffusion models
Diffusion Models (Sohl-Dickstein et al. 2015; Ho et al. 2020), the focus of this work, systematically corrupt data with noise and then learn to reverse the process. They have been shown to achieve superior sample quality and mode coverage compared to other classes of models (Dhariwal & Nichol 2021), making them a compelling choice for modelling the complex and diverse morphologies of galaxies observed by Euclid.
@@
This paper makes use of one such implementation, denoising diffusion probabilistic models (DDPMs, Ho et al. 2020). Diffusion works on the following premise: by continually adding a small amount of noise to an image, eventually, one is left with an image of complete noise without any remnants of the original input. If the noise is generated in a stochastic but consistent way, such as from a Gaussian distribution, we can use a network to learn the dynamics of the noise and perform a reverse process. By predicting the noise that was added to an image $x$ at timestep $t$, we can remove this noise to produce image $x_{t-1}$. This forms an iterative process originating from some final timestep $T$, an image of pure noise, to a realistic image from the data distribution at timestep 0.
The formal definition of this diffusion process, following Ho et al. (2020) and Nichol & Dhariwal (2021), is expressed as $q(x_{1},\dots,x_{T})$ which represents the joint distribution of a sequence where each image $x_{t}$ is progressively noised.
@@
where, $q(x_{t}|x_{t-1})$ specifies the conditional distribution at each timestep $t$, modelling the incremental addition of noise:
@@
With a large enough $T$ and an adjusted time-dependent variance $\beta_{t}$, $x_{T}$ will approach a Gaussian distribution that is isotropic, meaning uniform in all directions. For the denoising process, starting at a $x_{T}\sim\mathcal{N}(0,I)$, accurately modelling $q(x_{t-1}|x_{t})$ is not tractable, requiring a neural network to approximate it. An example of the full diffusion process can be seen in the top panel of Fig. 1.
Background noise versus diffusion noise
@@
It is important to clarify the distinctions between the definition and use of ‘noise’ in diffusion models compared to the traditional understanding of noise in astronomy-based data. In the context of astronomical imaging, noise refers to the random fluctuations inherent in observational data, often caused by e.g. background emission, electronics read-out noise, and atmospheric disturbances. This ‘background noise’ reduces the signal-to-noise ratio (S/N) of the image, affecting the ability to detect subtle details in any analysis.
@@
In contrast, the diffusion process uses an intentional integration of Gaussian noise to facilitate the training of the generative model. The pixels of image $x_{T}$, follow the distribution: $\mathcal{N}(0,I)$, rather than following the distribution of the background noise. The goal of the reverse diffusion process is to remove the artificial, high-variance Gaussian noise that was intentionally added during the forward process, thereby transforming a sample from a simple prior distribution ($\mathcal{N}(0,I)$) into a realistic galaxy image. A realistic galaxy image, as defined by our training data set, inherently includes the source’s light profile superimposed on the characteristic background noise of the Euclid instrument. Therefore, the model learns to generate images that replicate this entire distribution, including the statistical properties of the background noise. The diffusion pipeline does not aim to produce a ‘denoised’ image in the traditional astronomical sense (i.e., a background-subtracted image).
Use in Astronomy Applications
@@
Smith et al. (2022) also utilised the DDPM framework to generate a data set of realistic galaxies. The model is trained using the Photometry and Rotation curve OBservations from Extragalactic Surveys (PROBES) data set (Stone & Courteau 2019; Stone et al. 2021), a collection of large, well resolved objects that feature significant internal structure. They provide analysis of the quality of their generated images, showcasing the diffusion model’s ability to create visually realistic images which feature similar physical property distributions such as half-light radius and flux-space colour values. They also briefly explore the use of inpainting with their model to remove satellite trails.
@@
When researching diffusion models, papers utilising score-based models are often used and spoken of interchangeably, as they are different implementations of the same generative process (Smith et al. 2022). Rather than the fixed sequence of timesteps to denoise data, score-based methods apply stochastic differential equations (SDEs) to estimate the data distribution’s gradient (score). Working with SDEs and gradients allows for a continuous range of possible diffusion paths. It is for this reason that score-based models can be classed as the more general framework for diffusion-based generative models. The use of score-based generative models within astronomical applications includes galaxy image deconvolution (Adam et al. 2023; Spagnoletti et al. 2024), gravitation lensing analysis (Adam et al. 2022; Remy et al. 2023), and deblending (Sampson et al. 2024).

### v3

@@
Light emission from galaxies exhibit diverse brightness profiles, influenced by factors such as galaxy type, structural features, and interactions with other galaxies. Elliptical galaxies feature more uniform light distributions, while spiral and irregular galaxies have complex, varied light profiles due to their structural heterogeneity and star-forming activity. In addition, galaxies with active galactic nuclei (AGN) feature intense, concentrated emission from gas accretion around supermassive black holes, superimposed on regular galactic light, while quasi-stellar objects (QSOs) represent extreme cases in which AGN emissions dominate their host galaxies. The challenge of identifying AGN and QSOs has been discussed many times in the literature, often requiring multi-wavelength observations. This paper introduces a novel approach to identify AGN and QSOs from a single image. Diffusion models have recently been developed in the machine-learning literature to generate realistic-looking images of everyday objects. Utilising the spatial resolving power of the Euclid VIS images, we created a diffusion model trained on one million sources, without using any source pre-selection or labels. The model learns to reconstruct light distributions of normal galaxies, since the population is dominated by them. We conditioned the prediction of the central light distribution by masking the central few pixels of each source and reconstructed the light according to the diffusion model. We further used this prediction to identify sources that deviate from this profile by examining the reconstruction error of the few central pixels regenerated in each source’s core. Our approach, solely using VIS imaging, features high completeness compared to traditional methods of AGN and QSO selection, including optical, near-infrared, mid-infrared, and X-rays. Our study offers practical insights for refining diffusion models and broadening their applications throughout the Euclid survey area, underscoring the utility of this approach in diverse astronomical contexts beyond just AGN identification.
@@ 1 Introduction
Ever since the discovery that the mass or dynamics of a spheroid within a galaxy correlates with the mass of any detectable supermassive black hole (SMBH) at its centre (e.g. Dressler 1989; Kormendy & Richstone 1995; Magorrian et al. 1998; Ferrarese & Merritt 2000, and more recently, Sahu et al. 2019; Davis et al. 2018, 2019 and references therein), it has been believed that most – if not all – galaxies with a spheroidal (or pseudo-spheroidal) component contain a SMBH at their centre. If a galaxy’s emission contains evidence of its central SMBH being actively fuelled, it will emit non-thermal radiation, leading it to be classified as an active galactic nucleus (AGN). Depending upon the luminosity of that emission and its contrast with the stellar emission from the surrounding galaxy, it could have a classification ranging from a Seyfert galaxy (for the weakest emission) through to a quasi-stellar object or QSO (for the most luminous).
@@
For imaging in wavebands such as the optical and near-IR, the presence and detectability of emission from the area directly around a SMBH depends on multiple factors. This can include the rate of infall of material into the region around the SMBH, the amount of obscuration (absorption or scattering) towards that region, the contrast between the SMBH-related emission and that of the surrounding galaxy, the redshift, and the intrinsic compactness of the stellar distribution of the host galaxy. The growth and mass of a central SMBH is intimately related to the evolution of its host galaxy, not least through the action of energy, momentum, and radiative feedback from outflows originating from the area around the SMBH on the properties of the stellar populations and ISM (see Harrison & Ramos Almeida 2024 for a recent review). Consequently, identifying and characterising SMBHs, particularly during periods of active feedback, is central to the study of SMBH and galaxy evolution.
@@
A more precise classification of a source as an AGN is possible when a much wider range of multi-wavelength data is available. This can include spectroscopy and considering the much broader spectral energy distribution of the galaxy and/or the candidate AGN. However, it is still worth considering the efficient selection of AGN candidates from single-band imaging as a first step in identifying large samples of AGN for further study. In particular, with the advent of the uniform, deep, and high-spatial-resolution optical imaging dataset from the Euclid mission, covering a large fraction of the extragalactic sky, there is the potential to derive unprecedentedly large samples of AGN across a wide range of the (optical) luminosity and redshift parameter space with the use of an appropriate and sufficiently efficient technique. The technique can then subsequently be refined by feeding back the result of using other multi-wavelength data, allowing us to confirm or reject candidates. With sufficient refinement, this further assessment step eventually becomes either unnecessary, or at the very least, an efficient stage in compiling reliable samples. This then opens the door to the production of unprecedentedly large and reliable samples of optically selected AGN.
@@
Figure 1: Diffusion pipeline (top) that progressively adds noise to images, training the model to predict what noise was added from the previous step. Once trained and during inference, the model takes pure Gaussian noise as input and can iteratively remove the noise until a realistic galaxy image remains. Repeat inference runs will provide a different and unique galaxy from those it was trained on. The Repaint pipeline (bottom) takes the trained diffusion model and enables conditioning to allow parts of an existing image to be preserved by masking. At each denoising step, noise levels in the preserved pixels are adjusted to ensure they integrate correctly with the newly generated sections. After $T$ iterations, the output includes the retained pixels and newly generated areas, creating a different yet plausible final image.
@@
This work presents a technique that can achieve the above goal. In the following, we describe the use of diffusion models for generating samples of AGN candidates using Euclid VIS $I_{\scriptscriptstyle\rm E}$ band (Euclid Collaboration: Mellier et al., 2025; Euclid Collaboration: Cropper et al., 2025) imaging alone. Diffusion-based methods are machine-learning models that generate or reconstruct data by evolving a random distribution of pixels into a structured output over several steps. This technique outperforms traditional methods such as median pixel substitution or interpolation, which often struggle to recreate complex structures within images. Using these models for ‘inpainting’, whereby we adaptively recreate parts of the image, we can measure the discrepancies between the original and generated pixels. Examining these errors allows us to separate AGN, QSOs, and other sources from the broader galaxy population.
@@
Our method [^1] The code is publicly available here: https://github.com/grant-m-s/Astro-Repaint. identifies potential AGN by treating the task as an unsupervised anomaly detection problem. It begins with an image of a galaxy, potentially containing AGN emission at its centre, and masks out the central pixels that might be influenced by such emission. A diffusion model, trained on a large, unlabeled dataset of galaxies – most of which are normal, non-active – then predicts or ‘inpaints’ the masked region based on the learned prior distribution of typical galaxy morphologies. Because AGN are rare, there is no requirement to remove them from the training sample; the model is inherently biased towards reconstructing the central region as if it belonged to a typical, non-active galaxy. As a result, any significant discrepancy between the inpainted and actual central region suggests the presence of an AGN. This reconstruction error thus serves as a powerful anomaly score for identifying candidate AGN.
@@
While this paper details the fundamental technique applied to VIS imaging data, as this approach utilises machine learning, future improvements could incorporate findings from a detailed assessment that involves additional datasets. These could include multi-wavelength cross-matching or spectroscopic validation, from which we can ultimately reduce the need for further time-expensive refinement steps. Such enhancements would help to further improve the accuracy and reliability of the AGN candidate samples generated by the technique. Additionally, the technique presented here uses a standard treatment of diffusion models, which in principle can be significantly further optimised for the specific characteristics of the Euclid dataset, thereby improving further the efficiency of the processing of the Euclid data. Such work would be the subject of future papers.
@@ 2.1 Generative models
Generative models are a class of machine learning systems designed to learn an underlying data distribution, $p(x)$, from a finite set of independent and identically distributed (i.i.d.) samples, $\{x_{i}\}^{N}_{i=1}\sim p(x)$. The primary goal of these models is to develop a tractable method for generating new, synthetic samples that are statistically indistinguishable from the training data. Over the years, several distinct families of generative models have emerged, each with unique architectural and training paradigms. A key distinction lies in their optimisation objectives, which often seek to minimise a measure of divergence between the true data distribution, $p(x)$, and the model’s learned distribution, $p_{\theta}(x)$. The choice of divergence has significant implications for model behavior. Prominent examples include variational autoencoders (VAEs, Kingma & Welling, 2013), which learn a latent variable model of the data, and generative adversarial networks (GANs, Goodfellow et al., 2014, 2020), which involve a contest between a generator and a discriminator. In practice, GANs implicitly optimise an objective related to the reverse Kullback-Leibler (KL) divergence, $D_{\mathrm{KL}}[p_{\theta}(x)||p(x)]$, which exhibits a ‘mode-seeking’ property. This encourages the generation of high-fidelity samples but can lead to mode collapse, whereby the model learns only a subset of the data distribution (Thanh-Tung & Tran, 2020). In contrast, VAEs and diffusion models optimise objectives related to the forward KL divergence, $D_{\mathrm{KL}}[p(x)||p_{\theta}(x)]$, which is ‘mass-covering’. This property encourages the model to account for all modes present in the data, leading to greater sample diversity and more comprehensive coverage of the data distribution.
@@ 2.2 Diffusion models
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), the focus of this work, systematically corrupt data with noise and then learn to reverse the process. They have been shown to achieve superior sample quality and mode coverage compared to other classes of models (Dhariwal & Nichol, 2021), making them a compelling choice for modelling the complex and diverse morphologies of galaxies observed by Euclid.
@@
This paper makes use of one such implementation, denoising diffusion probabilistic models (DDPMs, Ho et al., 2020). Diffusion works on the following premise: by continually adding a small amount of noise to an image, eventually, one is left with an image of complete noise without any remnants of the original input. If the noise is generated in a stochastic but consistent way, such as from a Gaussian distribution, we can use a network to learn the dynamics of the noise and perform a reverse process. By predicting the noise that was added to an image, $x$, at timestep $t$, we can remove this noise to produce image $x_{t-1}$. This forms an iterative process originating from some final timestep, $T$, an image of pure noise, to a realistic image from the data distribution at timestep 0.
The formal definition of this diffusion process, following Ho et al. (2020) and Nichol & Dhariwal (2021), is expressed as $q(x_{1},\dots,x_{T})$, which represents the joint distribution of a sequence in which each image, $x_{t}$, is progressively noised.
@@
where $q(x_{t}|x_{t-1})$ specifies the conditional distribution at each timestep, $t$, modelling the incremental addition of noise:
@@
With a large enough $T$ and an adjusted time-dependent variance, $\beta_{t}$, $x_{T}$ will approach a Gaussian distribution that is isotropic, meaning uniform in all directions. For the denoising process, starting at a $x_{T}\sim\mathcal{N}(0,I)$, accurately modelling $q(x_{t-1}|x_{t})$ is not tractable, requiring a neural network to approximate it. An example of the full diffusion process can be seen in the top panel of Fig. 1.
2.2.1 Background noise versus diffusion noise
@@
It is important to clarify the distinctions between the definition and use of ‘noise’ in diffusion models compared to the traditional understanding of noise in astronomy-based data. In the context of astronomical imaging, noise refers to the random fluctuations inherent in observational data, often caused by, for example, background emission, electronics read-out noise, and atmospheric disturbances. This ‘background noise’ reduces the signal-to-noise ratio (S/N) of the image, affecting the ability to detect subtle details in any analysis.
@@
In contrast, the diffusion process uses an intentional integration of Gaussian noise to facilitate the training of the generative model. The pixels of image $x_{T}$ follow the distribution $\mathcal{N}(0,I)$, rather than following the distribution of the background noise. The goal of the reverse diffusion process is to remove the artificial, high-variance Gaussian noise that was intentionally added during the forward process, thereby transforming a sample from a simple prior distribution $\mathcal{N}(0,I)$] into a realistic galaxy image. A realistic galaxy image, as defined by our training dataset, inherently includes the source’s light profile superimposed on the characteristic background noise of the Euclid instrument. Therefore, the model learns to generate images that replicate this entire distribution, including the statistical properties of the background noise. The diffusion pipeline does not aim to produce a ‘denoised’ image in the traditional astronomical sense (i.e. a background-subtracted image).
2.2.2 Use in astronomy applications
@@
Smith et al. (2022) also utilised the DDPM framework to generate a dataset of realistic galaxies. The model was trained using the Photometry and Rotation curve OBservations from Extragalactic Surveys (PROBES) dataset (Stone & Courteau, 2019; Stone et al., 2021), a collection of large, well-resolved objects that feature significant internal structure. They provide analysis of the quality of their generated images, showcasing the diffusion model’s ability to create visually realistic images that feature similar physical property distributions such as the half-light radius and flux-space colour values. They also briefly explore the use of inpainting with their model to remove satellite trails.
@@
When researching diffusion models, papers utilising score-based models are often used and spoken of interchangeably, as they are different implementations of the same generative process (Smith et al., 2022). Rather than the fixed sequence of timesteps to denoise data, score-based methods apply stochastic differential equations (SDEs) to estimate the data distribution’s gradient (score). Working with SDEs and gradients allows for a continuous range of possible diffusion paths. It is for this reason that score-based models can be classed as the more general framework for diffusion-based generative models. The use of score-based generative models within astronomical applications includes galaxy image deconvolution (Adam et al., 2025; Spagnoletti et al., 2024), gravitation lensing analysis (Adam et al., 2022; Remy et al., 2023), and deblending (Sampson et al., 2024).